OpenStudy (anonymous):

Christie has a total of 15 pieces of fruit, all bananas and apples, worth $1.59. Bananas are 13 cents each and apples are 7 cents each. How many bananas and how many apples does she have? Select one: a. 6 bananas, 9 apples b. 9 bananas, 6 apples c. 9 bananas, 24 apples d. 21 bananas, 6 apples

OpenStudy (anonymous):

Let 'a' be the amount of apples and 'b' the amount of bananas a + b = 15 (she has 15 pieces of fruit total) 0.07a + 0.13b = 1.59 (the fruit cost $1.59 total) The first equation can be written b = 15 - a. Now substitute that into the second equation. 0.07a + 0.13(15 - a) = 1.59 0.07a + 1.95 - 0.13a = 1.59 -0.06a + 1.95 = 1.59 -0.06a = -.36 a = -0.36 / -0.06 a = 6 Now plugging that back into the first equation: b = 15 - 6 = 9. so apples= 6 bananas = 9